# On the observation of cosmic strings via gravitational-wave lensing

### v2

@@ Appendix C Detection rate
In this appendix, we explain the computation of the lensing detection rate for cosmic strings. We consider a network with three O3 detectors and impose three detection criteria. First, the unlensed waveform must be detectable with a large enough signal-to-noise ratio (in our case, we choose $SNR\geq 8$). Second, lensing-induced interference fringes must be statistically significant. We require the lensed and unlensed waveforms to raise a mismatch value of $0.05$ or larger. Third, the parameter $\Delta$ must satisfy the condition imposed in Eq. (26). These combined criteria ensure that both the signal and its lensing signature are detectable and distinguishable. Now, to compute the detection rate of lensed waveforms we must integrate over all possible source binaries and cosmic string configurations that satisfy the above criteria. This integral is given by [42]:
@@
where $n_{S}(M)$ is the merger rate density per year and per comoving volume as a function of the total mass, $M$, and $V(\Delta,M)$ is a six-dimensional volume in the space of string and source locations that yields detectable lensing. This $V$ accounts for the spatial configurations of one string and one source that produce a fringe meeting the criteria. Here $V(\Delta,M)$ has units of (comoving volume)2 and essentially counts the volume of all source-string configurations (for a given $M$ and $\Delta$) that would be observable as a lensed fringe event. This volume factorizes as:
$V(\Delta,M)\;=\;\int_{0}^{\chi_{\max}}4\pi\,\chi_{L}^{2}\;V_{S}(\chi_{L},\Delta,M)\;d\chi_{L},$ (52)
where $V_{S}(\chi_{L},\Delta,M)$ is the 3D volume of total mass $M$ source locations that produce detectable fringes when a string of tension $\Delta$ is at comoving distance $\chi_{L}$. We integrate $\chi_{L}$ from 0 out to the detector’s horizon, $\chi_{\max}(M)$. For a given string at $\chi_{L}$, several conditions limit $V_{s}$. First, the source must lie beyond the string (so that the string is between source and observer). If the source comoving distance is $\chi_{S}$, we require $\chi_{S}>\chi_{L}$. Second, the source must be within the detector’s horizon: $\chi_{S}\leq\chi_{\max}(M)$, where $\chi_{\max}(M)$ is the maximum detection comoving distance for a binary of mass $M$ given $SNR\geq 8$. This $\chi_{\max}$ is essentially set by the SNR threshold and corresponds to the unlensed detection range for that mass. Finally, the source’s alignment relative to the string must yield observable fringes. This alignment is quantified by the dimensionless impact parameter $y=\theta_{L}/\Delta$, where $\theta_{L}$ is the angular separation between the source and the line-of-sight passing through the string, measured at the string. Essentially, $y=0$ means the source is directly behind the string, and $y=1$ corresponds to the source lying just at the edge of the string’s gravitational influence (roughly at the deficit angle $\Delta$). We define $y_{\max}$ as the maximum impact parameter that still satisfies the fringe detectability conditions (mismatch and resolution). This $y_{\max}$ defines the maximum angular separation between the source and the line of sight measured at the observer, $\varphi_{\max}$:
@@
where $d_{LS}=\frac{\chi_{S}-\chi_{L}}{1+z_{S}}$ is the angular distance between lens and source [105]. Since the angle $\theta_{\max}=y_{\max}\Delta$ is very small (of the order of $\Delta$), we can approximate this as:
@@
where we used that $d_{S}=\frac{\chi_{S}}{1+z_{S}}$. Then, for given $\chi_{L}$ and $\chi_{S}$, the area of the sphere of radius $\chi_{S}$ containing the sources meeting the criteria can be approximated by $\varphi_{\max}\chi_{S}L(\chi_{L},\chi_{S})$, $L$ being the length of the geodesic joining the two points where the cosmic string cuts the sphere of radius $\chi_{S}$:
@@
By integrating for all distances between $\chi_{L}$ and $\chi_{\max}(M)$, we get the source volume for a given string at $\chi_{L}$:
$V_{S}(\chi_{L},\Delta,M)=\int_{\chi_{L}}^{\chi_{\max}(M)}\varphi_{\max}L(\chi_{L},\chi_{S})\,\chi_{S}\,d\chi_{S}$ (56)
with $\varphi_{\max}(\chi_{S},\chi_{L},M,\Delta)$ given by Eq. 54 and $L(\chi_{L},\chi_{S})$ by Eq. 55.
The overall detection rate is finally obtained by integrating over the mass distribution:
$R(\Delta)=\int dMn_{S}(M)n_{L}V_{6D}(\Delta,M)$ (57)
In practice, we integrate by summing across discrete mass and distance bins. We assume an optimistic astrophysical merger rate model for $n_{S}(M)$ (the M10 model) and a constant density $n_{L}=N_{L}/V_{H}$, where $N_{L}$ is the number of strings inside a Hubble volume $V_{H}$. We will take $N_{L}=30$, following the estimations done in [36]. The results of these calculations are shown in Figure 11.
@@
We observe that, for small values of $\Delta$, the rates are lower. The reason is that the condition imposed by Eq. (26) makes the closest lenses (small $\chi_{L}$) invisible.

### v3

@@ Appendix C Detection rate
In this appendix, we describe how we compute the lensing detection rate for cosmic strings. We consider a network of three O3 detectors and impose three detection criteria. First, the unlensed waveform must be detectable with a sufficiently large signal-to-noise ratio (in this work, we require $SNR\geq 8$). Second, the lensing-induced interference fringes must be statistically significant, which we enforce by requiring a mismatch of at least $0.05$ between the lensed and unlensed waveforms. Third, the parameter $\Delta$ must satisfy the condition given in Eq. (26). Taken together, these criteria ensure that both the signal and its lensing imprint are detectable and distinguishable. Now, to compute the detection rate of lensed waveforms, we integrate over all possible source binaries and cosmic string configurations that satisfy the above criteria. The resulting expression for the detection rate is given by [43]
@@
where $n_{L}$ is the density of cosmic strings per comoving volume, $n_{S}(M)$ is the merger rate density per year and per comoving volume as a function of the total mass of the binary system, $M=M_{1}+M_{2}$, and $V(\Delta,M)$ is a six-dimensional volume in the space of string and source locations that yields detectable lensing. This $V$ accounts for all spatial configurations of one string and one source that produce an interference fringe satisfying the detection criteria. Accordingly, $V(\Delta,M)$ has units of (comoving volume)2 and effectively counts all source-string configurations (for a given $M$ and $\Delta$) that would be observable as a lensed fringe event. This volume factorizes as
$V(\Delta,M)\;=\;\int_{\chi_{\min}}^{\chi_{\max}}4\pi\,\chi_{L}^{2}\;V_{S}(\chi_{L},\Delta,M)\;d\chi_{L},$ (54)
where $V_{S}(\chi_{L},\Delta,M)$ is the 3D volume of source locations, for binaries of total mass $M$, that produce detectable fringes when a string of tension $\Delta$ is located at comoving distance $\chi_{L}$. The integration over $\chi_{L}$ runs from the minimum comoving distance to the string $\chi_{\min}(\Delta)$ determined by Eq. (26), up to the detector’s horizon, $\chi_{\max}(M)$.
For a given string at $\chi_{L}$, several conditions limit $V_{S}$. First, the source must lie behind the string (so that the string is between the source and the observer). If $\chi_{S}$ denotes the source comoving distance, then this requires $\chi_{S}>\chi_{L}$. Second, the source must be within the detector horizon: $\chi_{S}\leq\chi_{\max}(M)$, where $\chi_{\max}(M)$ is the maximum comoving distance at which a binary of mass $M$ can be detected with $SNR\geq 8$. This $\chi_{\max}$ is essentially set by the SNR threshold and corresponds to the unlensed detection range for that mass. Finally, the relative alignment of the source with respect to the string must produce observable interference fringes. This alignment is quantified by the dimensionless impact parameter $y=\theta/\Delta$, where $\theta$ is the angular separation between the source and the line of sight passing through the string, evaluated at the string position (see also Sec. III.2, where the parameter $y$ is introduced in the expression for the transmission factor). In this parametrization, $y=0$ corresponds to a source located along the line of sight directly behind the string, while $y=1$ corresponds to a source lying at the boundary of the double-imaging region, set by the deficit angle $\Delta$, which approximately marks the extent of the string’s gravitational influence. We define $y_{\max}$ as the maximum impact parameter for which the fringe detectability conditions (mismatch and resolvability) are satisfied. This $y_{\max}$ then determines the maximum angular separation between the source and the line of sight as measured at the observer, $\varphi_{\max}$:
@@
where $d_{LS}=(\chi_{S}-\chi_{L})/(1+z_{S})$ is the angular diameter distance between the lens and the source [106]. Since the angle $\theta_{\max}=y_{\max}\,\Delta$ is very small (of the order of $\Delta$), we can approximate $\varphi_{\max}$ as
@@
where we have used $d_{S}=\chi_{S}/(1+z_{S})$. For given $\chi_{L}$ and $\chi_{S}$, the area of the sphere of radius $\chi_{S}$ containing the sources that satisfy the criteria can then be approximated by $\varphi_{\max}\,\chi_{S}\,L(\chi_{L},\chi_{S})$, where $L$ is the length of the geodesic connecting the two points at which the cosmic string intersects the sphere of radius $\chi_{S}$:
@@
By integrating over all source distances from $\chi_{L}$ up to the detector horizon $\chi_{\max}(M)$, we obtain the source volume corresponding to a string located at $\chi_{L}$:
$V_{S}(\chi_{L},\Delta,M)=\int_{\chi_{L}}^{\chi_{\max}}\varphi_{\max}\,L(\chi_{L},\chi_{S})\,\chi_{S}\,d\chi_{S}\,$, (58)
where $\varphi_{\max}(\chi_{S},\chi_{L},M,\Delta)$ is given by Eq. (56) and $L(\chi_{L},\chi_{S})$ by Eq. (57).
The overall detection rate is finally obtained by integrating Eq. (53) over the binary mass distribution:
$R(\Delta)$ $=\int dM\,n_{S}(M)\,n_{L}\,V(\Delta,M)$
$=8\pi n_{L}\Delta\,\int dM\,n_{S}(M)\int d\chi_{S}\int d\chi_{L}\,y_{\max}$
$\times\chi_{L}^{2}\,\chi_{S}^{2}\,\left(1-\frac{\chi_{L}}{\chi_{S}}\right)\,\arcsin\!\sqrt{1-\left(\frac{\chi_{L}}{\chi_{S}}\right)^{2}}.$ (59)
The dependence on the string tension $\Delta$ enters not only through the overall prefactor, but also implicitly via the integration limits on $\chi_{L}$ [see Eq. (54)]. In practice, the integrals are evaluated by summing over discrete bins in mass and distance. We adopt an optimistic astrophysical merger-rate model for $n_{S}(M)$ (the M10 model) and assume a constant cosmic-string density $n_{L}=N_{L}/V_{H}$, where $N_{L}$ denotes the number of strings within a Hubble volume $V_{H}$. Following Ref. [37], $N_{L}$ is estimated using the relation $n_{L}=N_{L}(z)H(z)^{3}$, which effectively counts the number of strings per Hubble volume. Using the parametrization developed in [107], Xiao et al. [37] find that, for an underlying phase transition occurring at $T\simeq 10^{15},\mathrm{GeV}$, the present-day network contains approximately $N_{L}(z=0)\simeq 30$ strings per Hubble volume. We adopt this value as our fiducial choice. Nevertheless, we note that other studies favor a smaller, order-unity number of long strings per Hubble volume, $N_{L}\sim\mathcal{O}(1)$ [108, 109]. Our constraints scale linearly with $N_{L}$: adopting a different number of strings per Hubble volume would rescale the predicted lensing rates by $N_{L}/30$ but would not qualitatively change our conclusions. The results of these calculations are shown in Fig. 11.
@@
We observe that the detection rates decrease for small values of $\Delta$. This behavior arises because the condition imposed by Eq. (26) excludes the closest lenses, corresponding to small values of $\chi_{L}$, from contributing to the rate.